OpenStudy (kj4uts):

1. Find sin(a - b) if sin(a)= 5/13 where a is in the second quadrant and cos(b)= -7/25 where b is in the third quadrant. 2. Find cos(a + b) if sin(a)= 8/17 where a is in the first quadrant and tan(b)= -7/24 where b is in the second quadrant. Please explain. Thank you!

OpenStudy (kj4uts):

This is what I got for number 1: sinα = 5/13 cos²α = 1-sin²α = 144/169 Since α is in the second quadrant, cosα = -√(144/169) = -12/13. cosβ = -7/25 sin²β = 1-cos²β = 576/625 Since β is in the third quadrant, sinβ = -√(576/625) = -24/25 I just do not know how to do the sum formula for sine? ...and I am not sure how to do number 2?

OpenStudy (faiqraees):

Are you not aware of the formula?

OpenStudy (faiqraees):

\[\large\rm sin(a-b) = sina* cosb - cos a *sinb \]\[\large\rm cos(a-b) = cosa* cosb + sin a *sinb \]

OpenStudy (kj4uts):

@FaiqRaees so like this sin(-12/13) * cos(-24/25) - cos(-24/25) * sin(-12/13)?

OpenStudy (faiqraees):

why have you changed the angle values?

OpenStudy (kj4uts):

What are the angle values?

OpenStudy (faiqraees):

sin(a)= 5/13 cos(b)= -7/25

OpenStudy (kj4uts):

Oh I see let me redo it then...

OpenStudy (kj4uts):

So like this sin(5/13) * cos(-7/25) - cos(5/13) * sin(-7/25)?

OpenStudy (kj4uts):

I do not see that as a choice?

OpenStudy (kj4uts):

That is the way of the formula?

OpenStudy (faiqraees):

Oh I see they havent given you values for sin b and cos a. So you have to find them out first

OpenStudy (faiqraees):

and then plug them in the equation

OpenStudy (faiqraees):

Can you do it?